LASER EXCITATION SPECTROSCOPY OF THE B̃^{2}A_{1} − X̃^{2}A_{1} TRANSITION OF CaOCH_{3}

C. J. WHITHAM AND J. M. BROWN, The Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford, OX1 3QZ, U.K.; \underline{S. A. BEATON}, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, British Columbia, Canada V6T 1Z1 ; Y. ITO, Department of Mechanical Engineering, Nagaoka University of Technology, 1603-1 Kamitomioka, Nagaoka, Niigata 940-21, Japan..

The B̃−X̃ transition of CaOCH_{3} has been observed at Doppler limited resolution and rotationally analysed. The molecule was generated using a laser ablation/molecular beam source in which a solid calcium rod was ablated and a mixture of argon seeded with a few percent of methanol was used as the carrier gas. The rotational analysis of the band is consistent with a linear Ca−O−C geometry in both electronic states and hence C_{3v} symmetry for the CaOCH_{3} radical. The values of the spin-rotation parameters determined for the excited B̃^{2}A_{1} state will be discussed.